# Distillation versus Contrastive Learning: How to Train Your Rerankers

### v1

@@ 5 Related Works
Text ranking with pre-trained transformers.
The application of transformers to text ranking has bifurcated into two primary architectural patterns: Bi-encoder retriever and cross-encoder reranker Humeau et al. (2020); Yates et al. (2021); Xu et al. (2025b). The synergy between these components forms the "retrieve and rerank" pipeline, which has become the standard paradigm for modern IR systems Asai et al. (2024); Singh et al. (2025); Xu et al. (2025b).
@@
In this paper, we focus exclusively on the architecture of cross-encoder pointwise reranker Nogueira and Cho (2019); Nogueira et al. (2020); Zhuang et al. (2024b). The reranker model processes a query and a document simultaneously, allowing the self-attention mechanism to explicitly model the interactions between their tokens throughout the layers. Nogueira and Cho (2019) first demonstrated the efficacy of this approach by fine-tuning BERT as a text pair classifier. Follow-up works are mainly investigating two directions: (1) addressing the token limitation to better handle the longer document text; (2) enhancing the contextualized representation used for the calculation of the ranking score. The fine-tuning of the BERT-type model as cross-encoder reranker has later been extended to encoder-decoder architectures, such as T5 Raffel et al. (2020); Nogueira et al. (2020), and decoder-only models like Llama Touvron et al. (2023); Ma et al. (2023).
@@
Irrespective of the model architecture, the fundamental training objective for a cross-encoder is to accurately discriminate between relevant and non-relevant documents for a given query. In practice, the non-relevant documents are usually hard negatives mined from first-stage retrievers (Nogueira et al., 2019; Gao et al., 2021; Boytsov et al., 2022, inter alia). Two primary training strategies are widely used to train cross-encoder reranker. The first is contrastive learning, which trains the model directly on ground-truth labels to maximize the distinction between positive and negative examples Gao et al. (2021); Zhuang et al. (2023). The second strategy is knowledge distillation, which involves training a student reranker to mimic the behavior and performance of a larger, more capable teacher model Hofstätter et al. (2020); Baldelli et al. (2024); Schlatt et al. (2025). In this work, we aim to compare these two training strategies in a controlled experimental setting to evaluate their strengths and weaknesses.
Knowledge distillation.
Knowledge distillation (KD) Hinton et al. (2015) facilitates the transfer of knowledge from a large, complex “teacher” model to a smaller, more efficient “student” model. The goal is to enable the student to mimic the teacher’s output, thereby inheriting its predictive capabilities at a reduced computational cost. Within the context of training neural IR models, knowledge distillation has been effectively applied to enhance both bi-encoder retriever (Hofstätter et al., 2020; Formal et al., 2021, inter alia) and cross-encoder reranker Baldelli et al. (2024); Schlatt et al. (2025).
TCT-ColBERT Lin et al. (2021b) distills the knowledge from a multi-vector representation, into a more computationally efficient single-vector representation that utilizes a simple dot product for similarity calculation. Hofstätter et al. (2020) demonstrated the efficacy of cross-architecture distillation, a method wherein knowledge from a cross-encoder reranker is distilled into a bi-encoder retriever. Their findings indicate that this approach can significantly improve the bi-encoder’s performance. Prior works have also explored knowledge distillation to train cross encoders Baldelli et al. (2024); Schlatt et al. (2025), but these explorations are limited to pointwise and pairwise loss functions. In contrast, listwise approaches train the student model to replicate the ranked order of a list of documents generated by the teacher for a given query. This is commonly achieved by minimizing the KL divergence between the probability distributions over the list of documents as predicted by the teacher and the student, thereby preserving the relative ranking Cao et al. (2007); Xia et al. (2008).
This work shows that training reranker by distilling knowledge from a larger teacher model generally yields better in-domain and out-of-domain ranking performance than contrastive learning based on ground-truth labels.
@@ 6 Conclusion and Future Work
In this paper, we aim to compare the effectiveness of two training strategies, i.e., contrastive learning and knowledge distillation in the context of training text rerankers. With a rigorous controlled experimental setup, we find that when distilling from a larger, more capable teacher model, rerankers trained with knowledge distillation achieve better in-domain and out-of-domain reranking performances compared to contrastive training strategy, and the observation is consistent across different model scales and language model architectures. Our future work will investigate a more optimized way to combine two training strategies, as well as to improve the robustness and out-of-domain generalization of knowledge distillation training.
@@ Appendix A Dataset Artifacts and Licenses
Appendix B Baseline Results
@@
We report the baseline results in Table 3.

### v2

@@ 5 Related Works
In this paper, we focus exclusively on the architecture of cross-encoder pointwise reranker Nogueira and Cho (2019); Nogueira et al. (2020); Zhuang et al. (2024b). The reranker model processes a query and a document simultaneously, allowing the self-attention mechanism to explicitly model the interactions between their tokens throughout the layers. Nogueira and Cho (2019) first demonstrated the efficacy of this approach by fine-tuning BERT as a text pair classifier. The fine-tuning of the BERT-type model as cross-encoder reranker has later been extended to encoder-decoder architectures, such as T5 Raffel et al. (2020); Nogueira et al. (2020), and decoder-only models like Llama Touvron et al. (2023); Ma et al. (2023).
@@
The training objective for a cross-encoder is to accurately discriminate between relevant and non-relevant documents for a given query. In practice, the non-relevant documents are usually hard negatives mined from first-stage retrievers (Nogueira et al., 2019; Gao et al., 2021; Boytsov et al., 2022, inter alia). Two primary training strategies are widely used to train cross-encoder reranker. The first is contrastive learning, which trains the model directly on ground-truth labels to maximize the distinction between positive and negative examples Gao et al. (2021); Zhuang et al. (2023). The second strategy is knowledge distillation, which involves training a student reranker to mimic the behavior and performance of a larger, more capable teacher model Hofstätter et al. (2020); Baldelli et al. (2024); Schlatt et al. (2025). Knowledge distillation (KD, Buciluǎ et al., 2006; Hinton et al., 2015) facilitates the transfer of knowledge from a large, complex “teacher” model to a smaller, more efficient “student” model. The goal is to enable the student to mimic the teacher’s output, thereby inheriting its predictive capabilities at a reduced computational cost. Within the context of training neural IR models, knowledge distillation has been effectively applied to enhance both bi-encoder retriever (Hofstätter et al., 2020; Formal et al., 2021, inter alia) and cross-encoder reranker Baldelli et al. (2024); Schlatt et al. (2025). In this work, we aim to compare these two training strategies in a controlled experimental setting to evaluate their strengths and weaknesses.
@@ 6 Conclusion and Future Work
In this paper, we compared the effectiveness of two training strategies, i.e., contrastive learning and knowledge distillation in the context of training text rerankers. With a rigorously controlled experimental setup, we find that when distilling from a larger, more capable teacher model, rerankers trained with knowledge distillation achieve better in-domain and out-of-domain reranking performances compared to contrastive training strategy, and the observation is consistent across different model scales and language model architectures. Our future work will investigate a more optimized way to combine two training strategies, as well as to improve the robustness and out-of-domain generalization of knowledge distillation training.
@@
This paper focuses on empirical experiments on public benchmarks. We believe this paper do not incur potential risks.
Acknowledgements
This material is based upon work supported in part by NSF DMS-2134223. This research is supported by the National Artificial Intelligence Research Resource (NAIRR) Pilot and the Delta advanced computing and data resource which is supported by the National Science Foundation (award NSF-OAC 2005572). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsers.
@@ Appendix A Dataset Artifacts and Licenses
Appendix B Hyperparameters
For all our training runs, we use a similar set of optimized hyperparameters identified from prior works and our preliminary experiments Ma et al. (2023); Xu et al. (2025a), only ablating learning rate to reduce the effect of overfitting. We use LoRA rank=16 and $\alpha$=32. We use AdamW optimizer, learning rate ranging from 3e-5 to 1e-4 with linear warmup and cool down. As we train all models on the train dataset for 1 epoch, we find learning rate is the most important hyperparameter to control overfitting. We use 8 GPUs with per device batch 4, gradient accumulation steps 4, which leads to a global batch size of 128. For each $(q_{i},d_{i}^{+})$ pair, we use $k=15$ negatives as we find increasing to 313131 negatives may lead to instable contrastive learning. For RepQwen training, we use in-batch negatives; while for reranker training, we focus solely on each $(q_{i},d_{i}^{+})$ pair’s own hard negatives.
Appendix C Baseline Results
@@
We report the baseline results in Table˜4.